a 6.90 molar sol of koh in water contains 30 % by weight of koh.find density of solution

Grade:12

1 Answers

sushant singh
13 years ago

hey richa, 

please check the solution...

to calculate the density of koh solution...

A  6.90 M solution contains 6.90 mol of KOH in 1000 cm3 of the solution.

Mass of 6.90 mol of KOH = 690 H 56 = 386.4 g 

A 30% solution contains 30g of KOH present in = 100 g of solution

∴ 386.4 g of KOH is present in =  (100 x 386.4 g) / 30 = 1288 g of solutions


Density of KOH solution = (Mass / volume) = 1288 / 1000 =1.288 g/cm-3